The area of a typical eardrum is about 5.00 x 10-
^5 m^2.
Calculate the sound power (the energy per second) incident on an eardrum at: a) the threshold of hearing, and, b) the threshold of pain
Answer #1

A = 5*10^-5

a)

β = 10 dB*log I/Io = 0

β = 0

I= Io= 1*10^-12 W/m2  

Power at whcih the threshold of hearing is

P = I*A = 1*10^-12* 5*10^-5 = 5*10^-17 W answer

B)

at the threshold level pain sound intensity  β  = 120 dB

I= Io*10^( β/10 )

= 10^-12*10^(120/10)
= 1 W/m^2

power P= I*A = 1*5*10^-5 = 5*10^-5 answer

Comment in case any doubt, will reply for sure.. Goodluck